# TMI! Finetuned Models Leak Private Information from their Pretraining Data

### v2

@@ 1 Introduction
Transfer learning has become an increasingly popular technique in machine learning as a way to leverage a model trained for one task to assist with building a model for a related task. Typically, we begin with a large pretrained model trained with abundant data and computation, and use it as a starting point for training a finetuned model to solve a new task where data and computation is scarce. This paradigm has been especially popular for privacy in machine learning [48, 64, 42, 12, 31, 29], because the data for pretraining is often considered public and thus the pretrained model provides a good starting point before we even have to touch sensitive data.
Although the data used to pretrain large models is typically scraped from the Web and publicly accessible, there are several reasons to believe that this data is still sensitive [57]. For example, personal data could have been published without consent by a third party who they trusted to keep their data private, and even ubiquitous and thoroughly examined pretraining datasets like ImageNet contain sensitive content [50, 61]. Beyond the privacy risks associated with individuals in the pretraining set, companies who utilize or sell finetuned models may also be at risk for privacy leakage. Consider the following example:
@@
Companies have large, web scraped datasets that are proprietary and remain internal (e.g., Google’s JFT-300 [60]). These datasets are used to train models that can be finetuned by individual teams within the company for their specific needs. These pretrained models are also hosted as a service where smaller companies can receive a finetuned model without ever seeing the pretrained model itself. For example, Google’s Vertex AI [1] allows smaller companies and individuals to upload their data and receive access to query the finetuned model as an endpoint. When these finetuned models are hosted publicly, they may leak sensitive information about the proprietary pretraining datasets on which they were trained.
Thus, a central question we attempt to understand in this work is: How much sensitive information does a finetuned model reveal about the data that was used for pretraining? We attempt to answer this question in both the settings where privacy preserving techniques have and have not been used to finetune the pretrained model. Examining these two settings leads to another research problem: Given that pretraining datasets have been shown to contain sensitive information [50, 61, 15], using the privacy preserving finetuning techniques described in prior work [48, 64, 42, 12, 31, 29] may not provide meaningful privacy guarantees in practice.
@@
Using the thought experiment from [57], consider a large, pretrained model, owned by Company A, that contains an individual’s sensitive data record. Suppose that this pretrained model is finetuned by Company B using differential privacy [23] with $(\varepsilon=0.5,\>\delta=10^{-5})$ on a sensitive downstream task. Considering that the open-source variants of these models’ pretraining datasets can exceed 5 TB (over 1 trillion tokens) in size [19], it is likely that any given individual’s data record can be present in both the pretraining and finetuning datasets. Thus, because differential privacy is not necessarily robust to preprocessing, the privacy guarantee from finetuning may not hold for individuals whose data record is present in both datasets.
@@
We study these questions via membership-inference (MI) attacks. A MI attack allows an adversary with access to the model to determine whether or not a given data point was included in the training data. These privacy attacks were first introduced by Homer et al. [34] in the context of genomic data, formalized and analyzed statistically by Sankararaman et al. [52] and Dwork et al. [24], and later applied in machine learning applications by Shokri et al. [53]. MI attacks have been extensively studied in machine learning applications, such as computer vision [13], contrastive learning [43], generative adversarial networks [17], and federated learning [47]. The success of MI attacks makes it clear that the pretrained model will leak information about the pretraining data. However, the process of finetuning the model will obscure information about the original model, and there are no works that study MI attacks that use the finetuned model to recover pretraining data.
@@
We comprehensively evaluate TMI on pretrained CIFAR-100 [39] and Tiny ImageNet [56] vision models, transferred to multiple downstream tasks. In our experiments with Tiny ImageNet, we evaluate the ability of TMI to infer membership on models finetuned on Caltech 101 [41]. Our pretrained CIFAR-100 models are finetuned on three downstream datasets of varying similarity to the pretraining data. In order of similarity to CIFAR-100, we evaluate TMI on models finetuned on a coarse-labeled version of CIFAR-100, CIFAR-10 [39], and the Oxford-IIT Pet dataset [49]. We also evaluate an extension of TMI on finetuned versions of publicly available large language language models, which are pretrained on WikiText-103 [46]. To measure the success of our attack we use several metrics, such as AUC and true positive rates at low false positive rates. To demonstrate the prevalence of privacy leakage with respect to pretraining data in finetuned models, we run TMI on target models with different finetuning strategies and settings with limited adversarial capabilities. We compare our results to both a simple adaptation of the likelihood ratio attack [13] to the transfer learning setting and a membership inference attack that has direct access to the pretrained model.
@@ 2.1.1 Scaling Model Confidences
Prior work [13] has used the logit function, $\texttt{logit}(p)=\smash{\log(\tfrac{p}{1-p})}$, to scale model confidences. This scaling yields an approximately normally distributed statistic that can be used to perform a variety of privacy attacks [13, 16, 58]. The logit function is obtained by inverting the sigmoid activation function, $\sigma(x)=\frac{1}{1+e^{x}}$, which is a specific case of softmax being used for binary classification.
@@ 2.1.3 Differential Privacy
Differential Privacy [23] is a mathematical definition of privacy that bounds the influence that any single individual in the training data has on the output of the model. Specifically, an algorithm satisfies differential privacy if for any two datasets that differ on one individual’s training data, the probability of seeing any set of potential models is roughly the same regardless of which dataset was used in training.
@@ 2.2.1 Privacy Attacks on Machine Learning Models
Deep learning models have been shown to memorize entire individual data points, even in settings where the data points have randomly assigned labels [66]. Prior work has demonstrated the ability of a wide class of deep learning models to perfectly fit training data while also achieving low generalization error [8]. In fact, recent work [27, 10, 28] has shown that memorization of training data may actually be necessary to achieve optimal generalization for deep learning models. As a result of this memorization, deep learning models tend to have higher prediction confidence on training data, which makes them highly susceptible to privacy attacks.
The most glaring violations of privacy in machine learning are reconstruction and training data extraction attacks. Early work in data privacy [22] showed that it is possible to reconstruct individuals’ data in statistical databases with access to noisy queries. More recently, training data extraction attacks have been shown to be successful when mounted on a variety of deep learning models, including large language models [15] and computer vision models [14].
Other attacks on machine learning models, such as membership-inference [53], property inference [7], and attribute inference [63] attacks are more subtle privacy violations. These attacks exploit vulnerabilities in machine learning models to learn whether or not an individual was in the training set, global properties of the training dataset, and an individual’s sensitive attributes. respectively. Recent versions of these attacks typically use a test statistic, such as loss [62] and model prediction confidences [63, 13, 16], to extract private information.
@@ 2.2.2 Membership-Inference Attacks
Membership-inference attacks [34] aim to determine whether or not a given individual’s data record was present in a machine learning model’s training dataset. These attacks represent a fundamental privacy violation that has a direct connection to differential privacy. Mounting these attacks and learning whether or not an individual was part of a sensitive dataset can serve as the basis for more powerful attacks. For example, prior work has used MI as a step in extracting training data [15]. Because of their simplicity, MI attacks are also a popular way to audit machine learning models for privacy leakage [54, 35, 62].
These attacks been extensively studied with two types of adversarial access: black-box query access and white-box access to the machine learning model’s parameters [40]. The query access setting has been more thoroughly studied, with attacks spanning several different machine learning domains, such as classification [13, 63, 53, 62, 58], natural language generation [13, 58], and federated learning [47]. Despite there being extensive work on black-box attacks and prior work on MI attacks on pretrained encoders [43], continuously updated models [38], and distilled models [36], there are few works that explore MI in the transfer learning setting where a pretrained model is finetuned on a new task. Zou et al. [69] study MI attacks that target individuals in the finetuning dataset, and Hidano et al. [33] explore ways in which an adversary can leverage control over the transfer learning process to amplify the success of MI attacks on the original model. No works have studied black-box MI attacks on the pretraining dataset of a finetuned machine learning model.
@@ 3 Threat Model
Our problem is to determine how much information a finetuned model reveals about the data used in the pretraining phase, and whether or not the finetuned model reveals strictly less information than the pretrained model. For this work we study this question using the language of MI attacks [34, 52, 24]. In the standard MI experiment and in our newly defined experiment, there is a machine learning model trained on some dataset, and a challenge point that is drawn from the same distribution as the training data. The challenge point is either an element of the training data or an independent point drawn from the same distribution. The attacker, who has access to the model and the challenge point, and knowledge of the distribution, tries to infer which of these two cases holds. In our experiment we separate the construction of the machine learning model into a pretraining phase and a finetuning phase, where the finetuning phase is performed with different data, drawn from a possibly different distribution. This finetuning phase introduces another layer of indirection that prevents the attacker from querying the original pretrained model, and thus potentially makes MI more difficult. Formally, our threat model, visualized in Figure 1, is described by the following game between a challenger $\mathcal{C}$ and an adversary $\mathcal{A}$:
@@
In our security game, we assume that the adversary has query access to the finetuned target model $f_{\beta}$ and knowledge of the pretraining data distribution $\mathbb{D}_{PT}$. Because we will be training shadow models [53] to perform our MI attack, the adversary also requires knowledge of the underlying distribution from which the finetuning dataset is sampled, $\mathbb{D}_{FT}$, and knowledge of the target model’s architecture and training algorithm. Although MI attacks vary in what they assume about the distribution and training algorithm [24], some degree of knowledge is necessary, so we consider full knowledge as a reasonable starting point. The knowledge we assume is the same as many other works on MI (e.g. [53, 13, 58, 63, 51, 36, 43]). We also assume that the adversary’s queries to the target model return numerical confidence scores for each label rather than just a single label, similar to prior privacy attacks [53, 63, 58, 13].
@@ 4.1 Membership Inference Under Distribution Shift
This specific choice of test statistic is motivated by prior work on membership-inference attacks on published statistics [25]. Subtracting the expectation of each term allows the adversary to observe whether the noise from computing $\hat{\mu}$ is correlated with the noise from sampling $c$. Thus, the test statistic $z$ is a real number that measures the correlation between the challenge point and the published statistic, $\hat{\mu}$. The adversary can then choose a threshold $\tau$ such that if $z>\tau$, they will predict that the challenge point was IN (i.e. $c\in X$). Else, the adversary will predict that the challenge point was OUT (i.e. $c\sim\mathcal{N}(\mu,\>\mathbb{I}_{d})$)
@@
[…]
@@
[…]
@@
Using Lemma 4.1, we can analyze the performance (AUC) of the adversary’s distinguishing test as a function of the parameter $\alpha$. To do this, we use the fact that the AUC of a classifier is equal to the probability that the classifier’s prediction on a randomly chosen positive (IN) sample is greater than the prediction on a randomly chosen negative (OUT) sample [26]. Here, we use the assumption that the test statistic is normal. Because $z$ is the inner product of two high dimensional Gaussian vectors, and thus the sum of many i.i.d. Gaussian random variables, as $d\to\infty$, $z$ is normally distributed.
@@
Figure 3 visualizes the attack’s AUC from Lemma 4.2 as a function of the parameter $\alpha$. Here, the parameters $n$, $m$, and $d$ are fixed. Our choices of the data’s dimension, $d$, are motivated by the dimension of the data in our evaluation on vision models (Section 5.3). We observe that there is a rapid increase in AUC as more information about $X$ is preserved in the public statistic (i.e. as $\alpha$ increases). In this figure, the value for $\alpha$ when $d=12,000$ and $\|\nu\|_{2}=5$ that minimizes the error on the new task, $Y$, is roughly 0.77. This $\alpha$ corresponds to an AUC of 0.96. If we make the distribution shift larger, say $\|\nu\|_{2}=10$, the optimal value of $\alpha$ is 0.51, which corresponds to an AUC of 0.78. This shows that the success of our attack on mean estimation depends on the extent to which we combine the means of $X$ and $Y$ using the parameter $\alpha$, which is based on the similarity of the "pretraining" data $X$ and the "finetuning" data $Y$.

### v3

@@ 1 Introduction
Transfer learning has become an increasingly popular technique in machine learning as a way to leverage a model trained for one task to assist with building a model for a related task. Typically, we begin with a large pretrained model trained with abundant data and computation, and use it as a starting point for training a finetuned model to solve a new task where data and computation is scarce. This paradigm has been especially popular for privacy in machine learning [papernot-2020-making, yu-2022-dpllm, li-2022-llmstrongdp, bu-2023-dpbiasft, he-2022-groupwisedp, ganesh-2023-pretraining], because the data for pretraining is often considered public and thus the pretrained model provides a good starting point before we even have to touch sensitive data.
Although the data used to pretrain large models is typically scraped from the Web and publicly accessible, there are several reasons to believe that this data is still sensitive [2022-tramer-position]. For example, personal data could have been published without consent by a third party who they trusted to keep their data private, and even ubiquitous and thoroughly examined pretraining datasets like ImageNet contain sensitive content [quach-2019-imagenet, yang-2021-imagenetfaces]. Beyond the privacy risks associated with individuals in the pretraining set, companies who utilize or sell finetuned models may also be at risk for privacy leakage. Consider the following example:
@@
Companies have large, web scraped datasets that are proprietary and remain internal (e.g., Google’s JFT-300 [JFT300]). These datasets are used to train models that can be finetuned by individual teams within the company for their specific needs. These pretrained models are also hosted as a service where smaller companies can receive a finetuned model without ever seeing the pretrained model itself. For example, Google’s Vertex AI [2023-vertexai] allows smaller companies and individuals to upload their data and receive access to query the finetuned model as an endpoint. When these finetuned models are hosted publicly, they may leak sensitive information about the proprietary pretraining datasets on which they were trained.
Thus, a central question we attempt to understand in this work is: How much sensitive information does a finetuned model reveal about the data that was used for pretraining? We attempt to answer this question in both the settings where privacy preserving techniques have and have not been used to finetune the pretrained model. Examining these two settings leads to another research problem: Given that pretraining datasets have been shown to contain sensitive information [quach-2019-imagenet, yang-2021-imagenetfaces, carlini-2020-extraction_llm], using the privacy preserving finetuning techniques described in prior work [papernot-2020-making, yu-2022-dpllm, li-2022-llmstrongdp, bu-2023-dpbiasft, he-2022-groupwisedp, ganesh-2023-pretraining] may not provide meaningful privacy guarantees in practice.
@@
Using the thought experiment from [2022-tramer-position], consider a large, pretrained model, owned by Company A, that contains an individual’s sensitive data record. Suppose that this pretrained model is finetuned by Company B using differential privacy [dwork-2006-dmns] with $(\varepsilon=0.5,\>\delta=10^{-5})$ on a sensitive downstream task. Considering that the open-source variants of these models’ pretraining datasets can exceed 5 TB (over 1 trillion tokens) in size [2023-llama-dataset], it is likely that any given individual’s data record can be present in both the pretraining and finetuning datasets. Thus, because differential privacy is not necessarily robust to preprocessing, the privacy guarantee from finetuning may not hold for individuals whose data record is present in both datasets.
@@
We study these questions via membership-inference (MI) attacks. A MI attack allows an adversary with access to the model to determine whether or not a given data point was included in the training data. These privacy attacks were first introduced by Homer et al. [homer-2008-mi] in the context of genomic data, formalized and analyzed statistically by Sankararaman et al. [SankararamanOJH09] and Dwork et al. [DworkSSUV15], and later applied in machine learning applications by Shokri et al. [shokri-2016-mi]. MI attacks have been extensively studied in machine learning applications, such as computer vision [carlini-2022-lira], contrastive learning [2021-liu-encodermi], generative adversarial networks [chen-2019-ganmi], and federated learning [nasr-2019-federatedmia]. The success of MI attacks makes it clear that the pretrained model will leak information about the pretraining data. However, the process of finetuning the model will obscure information about the original model, and there are no works that study MI attacks that use the finetuned model to recover pretraining data.
@@
We comprehensively evaluate TMI on pretrained CIFAR-100 [CIFAR] and Tiny ImageNet [tiny-imagenet] vision models, transferred to multiple downstream tasks. In our experiments with Tiny ImageNet, we evaluate the ability of TMI to infer membership on models finetuned on Caltech 101 [caltech-101]. Our pretrained CIFAR-100 models are finetuned on three downstream datasets of varying similarity to the pretraining data. In order of similarity to CIFAR-100, we evaluate TMI on models finetuned on a coarse-labeled version of CIFAR-100, CIFAR-10 [CIFAR], and the Oxford-IIT Pet dataset [pets]. We also evaluate an extension of TMI on finetuned versions of publicly available large language language models, which are pretrained on WikiText-103 [wikitext103]. To measure the success of our attack we use several metrics, such as AUC and true positive rates at low false positive rates. To demonstrate the prevalence of privacy leakage with respect to pretraining data in finetuned models, we run TMI on target models with different finetuning strategies and settings with limited adversarial capabilities. We compare our results to both a simple adaptation of the likelihood ratio attack [carlini-2022-lira] to the transfer learning setting and a membership inference attack that has direct access to the pretrained model.
@@ 2.1.1 Scaling Model Confidences
Prior work [carlini-2022-lira] has used the logit function, $\texttt{logit}(p)=\smash{\log(\tfrac{p}{1-p})}$, to scale model confidences. This scaling yields an approximately normally distributed statistic that can be used to perform a variety of privacy attacks [carlini-2022-lira, chaudhari-2022-snap, tramer-2022-truthserum]. The logit function is obtained by inverting the sigmoid activation function, $\sigma(x)=\frac{1}{1+e^{x}}$, which is a specific case of softmax being used for binary classification.
@@ 2.1.3 Differential Privacy
Differential Privacy [dwork-2006-dmns] is a mathematical definition of privacy that bounds the influence that any single individual in the training data has on the output of the model. Specifically, an algorithm satisfies differential privacy if for any two datasets that differ on one individual’s training data, the probability of seeing any set of potential models is roughly the same regardless of which dataset was used in training.
@@ 2.2.1 Privacy Attacks on Machine Learning Models
Deep learning models have been shown to memorize entire individual data points, even in settings where the data points have randomly assigned labels [zhang-2017-rethinking]. Prior work has demonstrated the ability of a wide class of deep learning models to perfectly fit training data while also achieving low generalization error [belkin-2019-reconciling]. In fact, recent work [feldman-2020-longtail, brown-2021-irrelevant, feldman-2020-memorize] has shown that memorization of training data may actually be necessary to achieve optimal generalization for deep learning models. As a result of this memorization, deep learning models tend to have higher prediction confidence on training data, which makes them highly susceptible to privacy attacks.
The most glaring violations of privacy in machine learning are reconstruction and training data extraction attacks. Early work in data privacy [dinur-2003-reconstruction] showed that it is possible to reconstruct individuals’ data in statistical databases with access to noisy queries. More recently, training data extraction attacks have been shown to be successful when mounted on a variety of deep learning models, including large language models [carlini-2020-extraction_llm] and computer vision models [carlini-2023-extraction_diffusion].
Other attacks on machine learning models, such as membership-inference [shokri-2016-mi], property inference [ateniese-2015-property_inference], and attribute inference [2018-yeom-mioverfitting] attacks are more subtle privacy violations. These attacks exploit vulnerabilities in machine learning models to learn whether or not an individual was in the training set, global properties of the training dataset, and an individual’s sensitive attributes. respectively. Recent versions of these attacks typically use a test statistic, such as loss [ye-2022-enhanced_mi] and model prediction confidences [2018-yeom-mioverfitting, carlini-2022-lira, chaudhari-2022-snap], to extract private information.
@@ 2.2.2 Membership-Inference Attacks
Membership-inference attacks [homer-2008-mi] aim to determine whether or not a given individual’s data record was present in a machine learning model’s training dataset. These attacks represent a fundamental privacy violation that has a direct connection to differential privacy. Mounting these attacks and learning whether or not an individual was part of a sensitive dataset can serve as the basis for more powerful attacks. For example, prior work has used MI as a step in extracting training data [carlini-2020-extraction_llm]. Because of their simplicity, MI attacks are also a popular way to audit machine learning models for privacy leakage [song-2018-auditingmi, tensorflow-2020-auditing, ye-2022-enhanced_mi].
These attacks been extensively studied with two types of adversarial access: black-box query access and white-box access to the machine learning model’s parameters [2020-klas-whitebox]. The query access setting has been more thoroughly studied, with attacks spanning several different machine learning domains, such as classification [carlini-2022-lira, 2018-yeom-mioverfitting, shokri-2016-mi, ye-2022-enhanced_mi, tramer-2022-truthserum], natural language generation [carlini-2022-lira, tramer-2022-truthserum], and federated learning [nasr-2019-federatedmia]. Despite there being extensive work on black-box attacks and prior work on MI attacks on pretrained encoders [2021-liu-encodermi], continuously updated models [jagielski-2023-updates], and distilled models [jagielski-2023-students], there are few works that explore MI in the transfer learning setting where a pretrained model is finetuned on a new task. Zou et al. [zhang-2020-transfer] study MI attacks that target individuals in the finetuning dataset, and Hidano et al. [hidano-2020-transmia] explore ways in which an adversary can leverage control over the transfer learning process to amplify the success of MI attacks on the original model. No works have studied black-box MI attacks on the pretraining dataset of a finetuned machine learning model.
@@ 3 Threat Model
Our problem is to determine how much information a finetuned model reveals about the data used in the pretraining phase, and whether or not the finetuned model reveals strictly less information than the pretrained model. For this work we study this question using the language of MI attacks [homer-2008-mi, SankararamanOJH09, DworkSSUV15]. In the standard MI experiment and in our newly defined experiment, there is a machine learning model trained on some dataset, and a challenge point that is drawn from the same distribution as the training data. The challenge point is either an element of the training data or an independent point drawn from the same distribution. The attacker, who has access to the model and the challenge point, and knowledge of the distribution, tries to infer which of these two cases holds. In our experiment we separate the construction of the machine learning model into a pretraining phase and a finetuning phase, where the finetuning phase is performed with different data, drawn from a possibly different distribution. This finetuning phase introduces another layer of indirection that prevents the attacker from querying the original pretrained model, and thus potentially makes MI more difficult. Formally, our threat model, visualized in Figure 1, is described by the following game between a challenger $\mathcal{C}$ and an adversary $\mathcal{A}$:
@@
In our security game, we assume that the adversary has query access to the finetuned target model $f_{\beta}$ and knowledge of the pretraining data distribution $\mathbb{D}_{PT}$. Because we will be training shadow models [shokri-2016-mi] to perform our MI attack, the adversary also requires knowledge of the underlying distribution from which the finetuning dataset is sampled, $\mathbb{D}_{FT}$, and knowledge of the target model’s architecture and training algorithm. Although MI attacks vary in what they assume about the distribution and training algorithm [DworkSSUV15], some degree of knowledge is necessary, so we consider full knowledge as a reasonable starting point. The knowledge we assume is the same as many other works on MI (e.g. [shokri-2016-mi, carlini-2022-lira, tramer-2022-truthserum, 2018-yeom-mioverfitting, sablayrolles-2019-whitebox, jagielski-2023-students, 2021-liu-encodermi]). We also assume that the adversary’s queries to the target model return numerical confidence scores for each label rather than just a single label, similar to prior privacy attacks [shokri-2016-mi, 2018-yeom-mioverfitting, tramer-2022-truthserum, carlini-2022-lira].
@@ 4.1 Membership Inference Under Distribution Shift
This specific choice of test statistic is motivated by prior work on membership-inference attacks on published statistics [dwork-2015-trace]. Subtracting the expectation of each term allows the adversary to observe whether the noise from computing $\hat{\mu}$ is correlated with the noise from sampling $c$. Thus, the test statistic $z$ is a real number that measures the correlation between the challenge point and the published statistic, $\hat{\mu}$. The adversary can then choose a threshold $\tau$ such that if $z>\tau$, they will predict that the challenge point was IN (i.e. $c\in X$). Else, the adversary will predict that the challenge point was OUT (i.e. $c\sim\mathcal{N}(\mu,\>\mathbb{I}_{d})$)
@@
${\mathbb{E}\mathopen{}\mathclose{{}\left(z}\right)}=0\quad\textrm{and}\quad{\mathrm{Var}\mathopen{}\mathclose{{}\left(z}\right)}=d\tilde{\alpha},$
@@
${\mathbb{E}\mathopen{}\mathclose{{}\left(z}\right)}=\frac{\alpha d}{n}\quad\textrm{and}\quad{\mathrm{Var}\mathopen{}\mathclose{{}\left(z}\right)}=d\tilde{\alpha}+\frac{2d\alpha^{2}}{n^{2}}$
@@
Using Lemma 4.1, we can analyze the performance (AUC) of the adversary’s distinguishing test as a function of the parameter $\alpha$. To do this, we use the fact that the AUC of a classifier is equal to the probability that the classifier’s prediction on a randomly chosen positive (IN) sample is greater than the prediction on a randomly chosen negative (OUT) sample [roc-analysis]. Here, we use the assumption that the test statistic is normal. Because $z$ is the inner product of two high dimensional Gaussian vectors, and thus the sum of many i.i.d. Gaussian random variables, as $d\to\infty$, $z$ is normally distributed.
@@
Figure 3 visualizes the attack’s AUC from Lemma 4.2 as a function of the parameter $\alpha$. Here, the parameters $n$, $m$, and $d$ are fixed. Our choices of the data’s dimension, $d$, are motivated by the dimension of the data in our evaluation on vision models (Section 5.3). We observe that there is a rapid increase in AUC as more information about $X$ is preserved in the public statistic (i.e. as $\alpha$ increases). In this figure, the value for $\alpha$ when $d=12,000$ and $\|\nu\|_{2}=5$ that minimizes the error on the new task, $Y$, is roughly 0.77. This $\alpha$ corresponds to an AUC of 0.96. If we make the distribution shift larger, say $\|\nu\|_{2}=10$, the optimal value of $\alpha$ is 0.51, which corresponds to an AUC of 0.78. This shows that the success of our attack on mean estimation depends on the extent to which we combine the means of $X$ and $Y$ using the parameter $\alpha$, which is based on the similarity of the ”pretraining” data $X$ and the ”finetuning” data $Y$.